Use Cramer’s rule to solve
x + 2y + z =5
2x + 2y + z =6
x + 2y + 3z =9 .
1
Expert's answer
2020-05-25T19:39:08-0400
 creating a matrix with coefficients in this equation 

"\begin{vmatrix}
 1 & 2& 1&5 \\
 2 & 2& 1& 6\\
1& 2&3 & 9
\end{vmatrix} 
(1) 






\begin{vmatrix}
 1 & 2& 1 \\
 2 & 2& 1\\
1& 2&3 
\end{vmatrix}
(2)" finding determinan of matrix (2) A = -4

Replace the 1st column of the main matrix with the solution vector(5,6,9) and find its determinant

"\begin{vmatrix}
 5 & 2& 1 \\
 6 & 2& 1\\
9& 2&3 
\end{vmatrix}" B = -4

Replace the 2nd column of the main matrix with the solution vector and find its determinant

"\begin{vmatrix}
 1 & 5& 1 \\
 2 & 6& 1\\
1& 9&3 
\end{vmatrix}" C = -4

Replace the 3rd column of the main matrix with the solution vector and find its determinant

"\begin{vmatrix}
 1 & 2& 5 \\
 2 & 2& 6\\
1& 2&9 
\end{vmatrix}" D = -8

x=B/A=1

y=C/A=1

z=D/A=2

Solution set:

x=1

y=1

z=2

check

1*1+2*1+1*2 = 5

2*1+2*1+1*2 = 6

1*1+2*1+3*2 = 9